A person is nearsighted and can clearly focus on objects that are no farther than 3 m away from her eyes. She borrows a friend's glasses but the borrowed glasses make things worse; that is, the person can now focus only on objects that are within 2.25 m away. What is the focal length of the borrowed glasses?
Answer #1

In this case image of the object at 2.25 m away is formed at 3 m away by the glass.

by lens formula,

\frac{1}{f} = \frac{1}{v} - \frac{1}{u}

= \frac{1}{-3} - \frac{1}{-2.25}

f = 9 m